OpenStudy (m_m13):

Combine the like terms to simplify the expression: 9 + 3 + 17xy(2) + 8y(3) + 10x –13xy(2) – 9x – 6y(3) A. 2y(3) + 4xy(2) + x + 12 B. 2y(3) + 4xy(2) + 10x + 9 C. 2y(3) – 4xy(2) – 3x + 12 D. 2y(6) + 4x2y(2) + 10x + 12 THE NUMBERS IN THE PARENTHESIS ARE EXPONENTS

OpenStudy (m_m13):

lease help I'm stuck

OpenStudy (texaschic101):

9 + 3 + 17xy^2 + 8y^3 + 10x - 13xy^2 - 9x - 6y^3 lets put our like terms together.. 8y^3 - 6y^3 = 2y^3 17xy^2 - 13xy^2 = 4xy^2 10x - 9x = x 9 + 3 = 12 so now we have : 2y^3 + 4xy^2 + x + 12

OpenStudy (m_m13):

thank you so much